Date:
Monday, August 24, 2015
Telescope:
CTIO Blanco 4-m
Submitter:
Michael Baumer
Submitter Email:
Observer Support:
Claudio Aguilera
Night Assistant:
None
Also Present:
Anna Schooneveld
Problems:
SISPI froze during a calibration sequence; configuration failed to remedy, therefore instance was successfully restarted..
Program:
Proposal:
2015B-0314
Target of Oportunity:
No
Time Observed:
0.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
5.5
Time lost other:
0.0
Description:
Catching Supernovae in the Act with KEGS (Kepler Extra-Galactic Survey) and DECam
Primary Investigator:
Brad Tucker
Email:
Institution:
U.Berkeley
Instrument:
CTIO 4m DECam
Organization:
DD
Total Program Time:
5.5
Proposal:
2012B-0001
Target of Oportunity:
No
Time Observed:
0.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
5.25
Time lost other:
0.0
Description:
Dark Energy Survey
Primary Investigator:
Josh Frieman
Email:
Institution:
Fermilab
Instrument:
CTIO 4m DECam
Organization:
NOAO
Total Program Time:
5.25
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-08-24 | 0.0 | 0.0 | 0.0 | 10.75 | 0.0 | 10.75 |
*Observing Team*
_DES Observers:_ Michael Baumer, Anna Schooneveld
_Telops:_ Claudio Aguilera
_Remote Support:_
*Observing Plan*
# Community observer first half: Director's discretionary
# Mid-night standards
# Turn on ObsTac
# Morning standards (while watching SISPI closely!)
*Conditions*
# Overcast, with rain in the vicinity. Dome remained closed for entire night.
*Accomplishments*
# successful PTC sequence (exposure 467111-477296)
*Narrative*
# Dome closed 1st half of night
# When conditions seemed to be statically bad, we decided to run a PTC sequence.
# 2 failed attempts: forgot to change proposal id and decided to add a flush_vsub, since it was unclear if this had been done, given the 1st half cancellation.
# 3rd attempt failed due to SISPI hang (see eLog) that required restart of SISPI instance.
# Finally successful from exposure 46711 onward.
# Dome remained closed for remainder of night.